OpenStudy (anonymous):

use the difference of squares rule to factor the following expression 9x^2-49y^2 What are they talking about??

OpenStudy (anonymous):

can u show me so I know

OpenStudy (2bornot2b):

\[(3x)^2-(7y)^2\implies (3x+7y)(3x-7y))\]

OpenStudy (hoblos):

a^2 - b^2 = (a-b)(a+b) you must use this formula

OpenStudy (2bornot2b):

They are talking about the difference of square formula i.e. \[a^2 - b^2 = (a-b)(a+b)\]

OpenStudy (anonymous):

Watch the video and become one with the force http://www.khanacademy.org/video/factoring-difference-of-squares?playlist=Developmental+Math+2

OpenStudy (anonymous):

Okay Thank you guys

OpenStudy (anonymous):

a^2 - b^2 = (a+b)(a-b)